OpenStudy (anonymous):

Which sequence can be represented by the expression 12n? 144, 132, 120, 108,… 1, 12, 24, 36,… 12, 24, 36, 48,… 6, 18, 30, 42,…

OpenStudy (anonymous):

clearly A 144 = 12x12 n = 12 132 = 12x11 n = 11 120 = 12x10 n = 10 108 = 12x9 n = 9

geerky42 (geerky42):

Seems like C works too.

OpenStudy (anonymous):

midpoint of AC=((-1+11) /2,( 5+13)/2)) i.e (5,9) midpoint of BD=((2+8)/2,(6+2)) i.e (5,4) slope of AC=(13-5)/(11-(-1))=8/12=2/3 Slope of BD=(6-2)/(2-8)=4/-6=-2/3 Slope of AC *Slope of BD=2/3*-2/3=-4/9 (not equal to -1) ABCD not a rhombus.